# HeightLane: BEV Heightmap guided 3D Lane Detection

### v1

@@ 1 Introduction
Figure 1: (a) Assuming the ground is a flat plane, 2D images or features can be transformed into BEV features using IPM. (b) Modeling the ground as a plane with 2 degrees of freedom (2-DoF), such as pitch and height, provides more generality and is used by LATR for positional encoding in the transformer. (c) Our method predicts a dense height map to spatially transform 2D image features onto a predefined BEV feature grid. Bold indicates how each method represents the ground.
Monocular 3D lane detection, which involves estimating the 3D coordinates of lane markings from a single image, is a fundamental task in autonomous driving systems. While LiDAR-based methods have achieved significant progress in many 3D perception tasks, monocular cameras are increasingly favored for 3D lane detection due to several key advantages. These advantages include lower costs, a superior perception range compared to LiDAR, and the ability to capture high-resolution images with detailed textures, which are essential for identifying narrow and elongated lane markings. Furthermore, the strong performance of deep learning-based 2D lane detection across various benchmarks has driven active research in this area, highlighting the potential for similar breakthroughs in 3D lane detection [27, 11, 13, 19, 29]. However, the lack of depth information in 2D images makes this task particularly challenging. Thus, accurately deriving 3D lane information from 2D images remains a significant research and development focus.
@@
To resolve the misalignment issues that arise from simplistic ground modeling, we propose HeightLane, a direct approach to ground modeling as shown in Fig. 1 (c). HeightLane creates a predefined BEV grid for the ground and generates multiple heightmap anchors on this grid, assuming various slopes. These heightmap anchors are then projected back onto the image to sample front-view features, allowing the model to efficiently predict a heightmap. To better align each BEV grid pixel with the 2D front-view features, height information from the predicted heightmap is added to the positional encoding of the BEV grid queries. Using the predicted heightmap along with deformable attention mechanisms, HeightLane explicitly performs spatial transformations of image features onto the BEV grid. This method significantly reduces the misalignment between the image and BEV features, ensuring more accurate representation and processing. By leveraging the heightmap for precise ground modeling, HeightLane effectively transforms front-view features into BEV features, thereby improving the accuracy and robustness of 3D lane detection.
@@ 3.2 Height guided Spatial Transform Framework
In this section, we propose a spatial transform framework utilizing the heightmap predicted in Sec. 3.1 as illustrated in Fig. 4. The BEV initial query is flattened and undergoes self-attention. During self-attention, BEV queries interact with each other, and positional encoding is added to each BEV query to provide positional information. The positional encoding is a learnable parameter. While studies performing attention on 2D FV features [14, 17] concatenate 3D ray coordinates with image feature queries, our method uses BEV grid coordinates and height embeddings for each BEV query. After the self-attention module, the output query of the self-attention module $\mathbf{Q}^{l}_{SA}$ in the $l^{th}$ layer is represented as follows:
@@ 3.3 Training
The loss between the predicted heightmap $\mathcal{H}$ and the ground truth heightmap $\mathcal{H}^{GT}$ is calculated using L1 loss.
$\mathcal{L}_{h}=\sum_{i=1}^{H^{\prime}}\sum_{j=1}^{W^{\prime}}\left|\mathcal{H}^{GT}_{ij}-\mathcal{H}_{ij}\right|$ (11)
@@ 4.1 Dataset
We evaluated our method using the OpenLane dataset [2], which encompasses a variety of road conditions, weather conditions, and lighting scenarios. OpenLane is built on the Waymo dataset [21], utilizing 150,000 images for training and 40,000 images for testing. The OpenLane dataset consists of 798 scenes for training and 202 scenes for validation, with each scene comprising approximately 200 images. Although OpenLane does not contain the information required to create heightmaps, it is based on Waymo, which allows us to extract the necessary LiDAR data from Waymo for each OpenLane scene. When extracting LiDAR data, we found that it is densely accumulated in the middle of each segment and becomes sparse towards the end frames. For example, Fig. 3 illustrates a scene where the car starts, goes uphill, turns right, and continues on another slope. At the starting point (green region), the LiDAR data is sparse, so bilinear interpolation was used to fill gaps in the heightmaps, ensuring consistency of the heightmap. The evaluation covers diverse scenarios, including Up & Down, Curve, Extreme Weather, Night, Intersection, and Merge & Split conditions. The evaluation metrics, as proposed by PersFormer[2], include the F-score, X-error, and Z-error for both near and far regions.
@@
HeightLane (Ours) 62.5 0.246 0.294 0.111 0.176
Table 2: Quantitative results comparison with other models on the OpenLane validation set. Evaluation metrics follow PersFormer[2]: F-score (higher is better), X-error, and Z-error (both lower are better). The best results are highlighted in bold and second-best results are underlined.
@@ 4.2 Implementation Details
We adopted ResNet-50[5] as the 2D backbone for extracting image features and set the image size to 576 x 1024. To obtain multi-scale image features, we added additional CNN layers to produce image features at 1/16 and 1/32 of the input image size, with each feature having 1024 channels. The BEV grid size for the heightmap and BEV feature was set to 200 x 48, with a resolution of 0.5 meters per pixel.
@@ 4.3.1 Qualitative Result
Fig. 5 shows a qualitative evaluation on the validation set of OpenLane. The predictions of the proposed HeightLane, the existing SOTA model LATR [17], and the ground truth are visualized. The ground truth is visualized in red, HeightLane in green, and LATR in blue. The first row of Fig. 5 displays the input images to the model, while the second row visualizes HeightLane, LATR, and the ground truth in 3D space. The third and fourth rows present the 3D lanes viewed from the Y-Z plane, where the Y-axis represents the longitudinal direction and the Z-axis represents height. The third row shows HeightLane and the ground truth, and the fourth row shows LATR and the ground truth.
Notably, HeightLane accurately detects lanes even in scenarios where the lanes are interrupted and resume, such as at intersections or over speed bumps. This is particularly evident in columns 1, 2, 4, 5, and 6 of the Fig. 5. In column 1, despite the occlusion from a car and partial lane markings, HeightLane continues to deliver precise lane predictions, demonstrating its robustness in handling complex scenes with occlusions and incomplete information. Additionally, thanks to the use of the heightmap, HeightLane effectively models changes in slope, as seen in column 3, where the road transitions from flat to sloped. In columns 2 and 5, which depict curved roads and partially visible lanes, HeightLane demonstrates superior prediction accuracy and maintains continuous lane detection even on curves, showcasing impressive performance.
@@ 4.3.2 Quantitative Result
Tab. 1 presents the quantitative evaluation of HeightLane. HeightLane achieved an overall F-score of 62.5% on the OpenLane validation set, outperforming all existing SOTA models. Specifically, HeightLane showed significant improvement in Extreme Weather, Night, and Intersection scenarios, achieving the best scores in these challenging conditions. Additionally, HeightLane demonstrated strong performance in Curve and Merge & Split scenarios, securing the second-best performance in these categories. Although HeightLane did not perform as well in the Up&Down scenario, this is because a 2-DoF plane assumption suffices for consistent uphill or downhill slopes. However, HeightLane excels in scenarios where the slope changes as column 3 in Fig. 5, demonstrating its capability to fit and predict more freely in varying gradient conditions.
Tab. 2 shows the F-score, X-error, and Z-error on the Openlane validation set. HeightLane achieved SoTA in terms of F-score, outperforming all other models with a score of 62.5%. Although it did not match the best-performing and second-best performing models in Z-error, it still demonstrated competitive results. In terms of X-error, HeightLane achieved the second-best performance, showcasing its robustness in estimating lane positions accurately in the lateral direction.
@@ 4.4 Ablation Study
Single-slope Heightmap Anchor 56.2
Multi-slope Heightmap Anchor 62.5
@@
Different Height Extraction Methods Tab. 3 shows the F-score corresponding to different height extraction methods. The view relation module, initially proposed in [18], is an MLP module used for transforming BEV features in [22]. The single-slope heightmap anchor method projects a zero-height plane onto the image and uses the sampled image features from this plane as the BEV features. However, this approach assumes a flat plane and samples only the 2D image features at this height, resulting in incomplete feature representation. Consequently, features corresponding to inclined roads are not sampled, even though they are part of the actual road. In contrast, the multi-slope heightmap anchor proposed in this paper projects multiple planes with various slopes onto the image, samples the image features from each plane, and fuses them to form the BEV features. This multi-anchor approach achieved the highest F-score.
@@
Heightmap Anchor Design Tab. 4 shows the F-scores for various heightmap anchor designs. The first row corresponds to the single-slope heightmap anchor from Tab. 3. When using 0° along with $\pm$3°, the performance improved by 4.5% compared to using only 0°. Similarly, using 0° along with $\pm$5° resulted in a 6.3% increase. Although the configuration with 0°, ±3°, and ±5° achieved the best performance, the difference was marginal compared to using just 0° and ±5°. However, increasing the number of heightmap anchors directly increases the number of channels in the final BEV height feature, which in turn raises the computational cost. To maintain a balance between performance and computational efficiency, we ultimately chose the configuration with 0° and ±5° heightmap anchors for the final method in the paper.
@@
Table 5: Comparison with multi-modal models on the OpenLane validation set. Ours (GT) means that we use ground truth heightmap for spatial feature transform framework during the inference step. M denotes the input modalities used, where M stands for camera and L stands for LiDAR.
Comparison with Multi-modal Methods Tab. 5 compares our method with various multi-modal 3D lane detectors. In this table, Ours (GT) represents the results obtained by using the ground truth heightmap instead of the height extraction module. This substitution aims to observe the performance of the spatial feature transform framework, assuming that the predicted heightmap from the height extraction module is highly accurate. By using the GT heightmap, which is derived from LiDAR data, we can make a fair comparison with detectors that utilize LiDAR input. The results suggest that when the heightmap is accurately predicted, our HeightLane method can achieve performance that is comparable to or even surpasses models employing both LiDAR and camera inputs. This demonstrates the potential of our approach to effectively utilize precise height information, highlighting the robustness and capability of our method in transforming between front-view features and BEV features.

### v2

@@ 1 Introduction
Figure 1: (a) Assuming the ground is a flat plane, 2D images or features can be transformed into BEV features using IPM[2]. (b) Modeling the ground as a plane with 2 degrees of freedom (2-DoF), such as pitch and height, provides more generality and is used by LATR [17] for positional encoding in the transformer. (c) Our method predicts a dense height map to spatially transform 2D image features onto a predefined BEV feature grid. Bold indicates how each method represents the ground.
Monocular 3D lane detection, which involves estimating the 3D coordinates of lane markings from a single image, is a fundamental task in autonomous driving systems. While LiDAR-based methods have achieved significant progress in many 3D perception tasks, monocular cameras are increasingly favored for 3D lane detection due to several key advantages. These advantages include lower hardware costs, a superior perception range compared to LiDAR, and the ability to capture high-resolution images with detailed textures, which are essential for identifying narrow and elongated lane markings. Furthermore, the strong performance of deep learning-based 2D lane detection across various benchmarks has driven active research in this area, highlighting the potential for similar breakthroughs in 3D lane detection [27, 11, 13, 19, 29]. However, the lack of depth information in 2D images makes this task particularly challenging. Thus, accurately deriving 3D lane information from 2D images remains a significant research and development focus.
@@
To resolve the misalignment issues that arise from simplistic ground modeling, we propose HeightLane, a direct approach to ground modeling as shown in Fig. 1 (c). HeightLane creates a predefined BEV grid for the ground and generates multiple heightmap anchors on this grid, assuming various slopes. These anchors are projected back onto the image to sample front-view features from the corresponding regions, enabling the model to efficiently predict a heightmap. To better align each BEV grid pixel with the 2D front-view features, height information from the predicted heightmap is added to the positional encoding of the BEV grid queries. Using the predicted heightmap along with deformable attention mechanisms, HeightLane explicitly performs spatial transformations of image features onto the BEV grid. This method significantly reduces the misalignment between the image and BEV features, ensuring more accurate representation and processing. By leveraging the heightmap for precise ground modeling, HeightLane effectively transforms front-view features into BEV features, thereby improving the accuracy and robustness of 3D lane detection.
@@ 3.2 Height guided Spatial Transform Framework
In this section, we propose a spatial transform framework utilizing the heightmap predicted in Sec. 3.1 as illustrated in Fig. 4. The BEV initial query is flattened and undergoes self-attention. During self-attention, BEV queries interact with each other, and positional encoding is added to each BEV query to provide positional information. The positional encoding is a learnable parameter. While studies performing attention on 2D front-view features [14, 17] concatenate 3D ray coordinates with image feature queries, our method uses BEV grid coordinates and height embeddings for each BEV query. After the self-attention module, the output query of the self-attention module $\mathbf{Q}^{l}_{SA}$ in the $l^{th}$ layer is represented as follows:
@@ 3.3 Training
The loss between the predicted heightmap $\mathcal{H}$ and the ground truth heightmap $\mathcal{H}^{GT}$ is calculated using Smooth L1 loss.
$\mathcal{L}_{\text{h}}=\begin{cases}\frac{1}{2}(\mathcal{H}^{GT}_{ij}-\mathcal{H}_{ij})^{2},&\text{if }|\mathcal{H}^{GT}_{ij}-\mathcal{H}_{ij}|<\beta,\\ |\mathcal{H}^{GT}_{ij}-\mathcal{H}_{ij}|-0.5,&\text{otherwise}.\end{cases}$ (11)
@@ 4.1 Dataset
We evaluated our method using the OpenLane dataset [2], which encompasses a variety of road conditions, weather conditions, and lighting scenarios. OpenLane is built on the Waymo dataset [21], utilizing 150,000 images for training and 40,000 images for testing. The OpenLane dataset consists of 798 scenes for training and 202 scenes for validation, with each scene comprising approximately 200 images. Although OpenLane does not contain the information required to create heightmaps, it is based on Waymo, which allows us to extract the necessary LiDAR data from Waymo for each OpenLane scene. When extracting LiDAR data, we found that it is densely accumulated in the middle of each segment and becomes sparse towards the end frames. For example, Fig. 3 illustrates a scene where the ego vehicle goes uphill, turns right, and continues on another slope. At the starting point (green region), the LiDAR data is sparse, so bilinear interpolation was used to fill gaps in the heightmaps, ensuring consistency of the heightmap. The evaluation covers diverse scenarios, including Up & Down, Curve, Extreme Weather, Night, Intersection, and Merge & Split conditions. The evaluation metrics, as proposed by PersFormer[2], include the F-score, X-error, and Z-error for both near and far regions.
@@
HeightLane (Ours) 62.7 0.240 0.266 0.116 0.165
Table 2: Quantitative results comparison with other models on the OpenLane validation set. The best results are highlighted in bold and second-best results are underlined.
@@ 4.2 Implementation Details
We adopted ResNet-50[5] as the 2D backbone for extracting image features and set the image size to 600 x 800. To obtain multi-scale image features, we added additional CNN layers to produce image features at 1/16 and 1/32 of the input image size, with each feature having 1024 channels. The BEV grid size for the heightmap and BEV feature was set to 200 x 48, with a resolution of 0.5 meters per pixel.
@@ 4.3.1 Qualitative Result
Fig. 5 shows a qualitative evaluation on the validation set of OpenLane. The predictions of the proposed HeightLane, the existing SOTA model LATR [17], and the ground truth are visualized. The ground truth is visualized in red, HeightLane in green, and LATR in blue. The first row of Fig. 5 shows the input images to the model. The second row visualizes HeightLane, LATR, and the ground truth in 3D space. The third and fourth rows display 3D lanes from the Y-Z plane, where the Y-axis represents the forward direction and the Z-axis represents height. The third row compares HeightLane to the ground truth, while the fourth compares LATR to the ground truth.
Notably, HeightLane accurately detects lanes even in scenarios where the lanes are interrupted and resume, such as at intersections or over speed bumps. This is particularly evident in columns 1, 2, 4, 5, and 6 of the Fig. 5. In column 1, despite the occlusion from a car and partial lane markings, HeightLane continues to deliver precise lane predictions, demonstrating its robustness in handling complex scenes with occlusions and incomplete information. Additionally, thanks to the use of the heightmap, HeightLane effectively models changes in slope, as seen in column 3, where the road transitions from flat to sloped. In columns 2 and 5, which depict curved roads and partially visible lanes, HeightLane demonstrates superior prediction accuracy and maintains continuous lane detection even on curves.
@@ 4.3.2 Quantitative Result
The evaluation metrics for quantitative assessment include the F-score, x error, and z error proposed by [2]. GT and predictions are matched based on the Euclidean distance, and a lane is classified as a true positive prediction depending on the proportion of matching points within the lane. Additionally, x and z errors are categorized into close-range (first 40 points) and far-range (remaining 60 points).
Tab. 1 presents the quantitative evaluation of HeightLane. HeightLane achieved an overall F-score of 62.7% on the OpenLane validation set, outperforming all existing SOTA models. Specifically, HeightLane showed significant improvement in Curve and Intersection scenarios, achieving the best scores in these challenging conditions. Additionally, HeightLane demonstrated strong performance in Up&Down and Merge&Split scenarios, securing the second-best performance in these categories. Although HeightLane did not achieve the highest score in the Up&Down scenario, it excelled in scenarios with changing slopes (column 3, Fig. 5), demonstrating its adaptability to varying gradient conditions.
Tab. 2 shows the F-score, X-error, and Z-error on the Openlane validation set. Although it did not match the best-performing and second-best performing models in Z-error, it still demonstrated competitive results. In terms of X-error, HeightLane achieved the second-best performance, showcasing its robustness in estimating lane positions accurately in the lateral direction.
@@ 4.4 Ablation Study
Single-slope Heightmap Anchor 57.1
Multi-slope Heightmap Anchor 62.7
@@
Different Height Extraction Methods Tab. 3 shows the F-score corresponding to different height extraction methods. The view relation module, initially proposed in [18], is an MLP module used for transforming BEV features in [22]. The single-slope heightmap anchor method projects a zero-height plane onto the image and uses the sampled image features from this plane as the BEV features. This approach assumes a flat plane, sampling only 2D image features at a fixed height, which leads to incomplete feature representation and excludes features of inclined or declined road. In contrast, the multi-slope heightmap anchor proposed in this paper projects multiple planes with various slopes onto the image, samples the image features from each plane, and fuses them to form the BEV features. This multi-anchor approach achieved the highest F-score.
@@
Heightmap Anchor Design Tab. 4 shows the F-scores for various heightmap anchor designs. Using 0° with $\pm$ 3° improved performance by 3.6%, while using 0° with $\pm$5° resulted in a 5.8% increase. Although the configuration with 0°, ±3°, and ±5° achieved the best performance, the difference was marginal compared to using just 0° and ±5°. Increasing the number of heightmap anchors raises the channels in the final BEV height feature and computational cost, so we balanced performance and efficiency by selecting 0° and ±5° anchors for the final method.
@@
Table 5: Comparison with multi-modal models on the OpenLane validation set. Ours (GT) means that we use ground truth heightmap for spatial feature transform framework. M indicates input modalities: C for camera and L for LiDAR.
Comparison with Multi-modal Methods Tab. 5 compares our method with various multi-modal 3D lane detectors. In this table, Ours (GT) represents the results obtained by using the ground truth heightmap instead of the height extraction module. This substitution aims to observe the performance of the spatial feature transform framework, assuming that the predicted heightmap from the height extraction module is highly accurate. By using the GT heightmap, which is derived from LiDAR data, we can make a fair comparison with detectors that utilize LiDAR input. The results show that accurate heightmap predictions enable HeightLane to match or surpass models using both LiDAR and camera inputs, highlighting its robustness in leveraging height information and transforming front-view to BEV features.
@@ 5 Conclusion
6 Acknowledgments
This work was partly supported by the Technology Innovation Program (No. 20018110, "Development of a wireless teleoperable relief robot for detecting searching and responding in narrow space") funded by the Ministry of Trade, Industry & Energy (MOTIE) and Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [No. RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)].